OpenStudy (anonymous):

Solve the equation |3m + 3 | = 15. m = 2 or m = −5 m = 4 or m = −6 m = 2 or m = 5 m = 4 or m = 6

hartnn (hartnn):

when \(|a| = b\) then we get a= b OR a= -b

hartnn (hartnn):

so what can you say about |3m+3| =15 ?

hartnn (hartnn):

let me give a similar example : |4x+9| =10 then we say 4x+9 =10 OR 4x+9 = -10